OpenStudy (anonymous):

What is the difference in volume between a 13 ft. cube and a 10 ft. cube? A. 1500 cubic ft. B. 1197 cubic ft. C. 3 cubic ft. D. 1300 cubic ft.

OpenStudy (campbell_st):

well calculate them out... the formula for volume of a cube is \[V = l^3\] where l is the side length... so you have \[V = 13^3~~~and ~~~V = 10^3\] calculate the values then find the difference.